Present Value of Annuity of $1 Periods 1.0% 2.0% 3.0% 3.75% 4.0% 4.25% 5.0% 6.0% 7.0% 0.99010 1.97040 2.94099 3.90197 4.85343Future Value of $1 Periods 1.0% 2.0% 3.0% 3.75% 4.0% 4.25% 5.0% 6.0% 7.0% 1.01000 1.02010 1.03030 1.04060 1.05101 1.06152 1.0Present Value of $1 Periods 1.0% 2.0% 3.0% 3.75% 4.0% 4.25% 5.0% 6.0% 7.0% 0.99010 0.98030 0.97059 0.96098 0.95147 0.94205 0.Future Value of Annuity of $1 Periods 1.0% 2.0% 3.0% 3.75% 4.0% 4.25% 5.0% 6.0% 7.0% 1.00000 2.01000 3.03010 4.06040 5.10101James Corporation is planning to issue bonds with a face value of $508,500 and a coupon rate of 6 percent. The bonds mature i
Answer #1

Solution:

a)

Case A – Issue Price – Market interest rate (annual): 4 percent

Face value of bonds = $508,500

Half yearly maturity period = 7 years * 2 = 14

Semiannual market interest = 4 /2 = 2%

Semiannual Cash Interest = Face value 508,500 * Coupon Rate 6% * ½ = $15,255

Issue Price = Cash Interest * Present value annuity factor at 2% for 14 period + Face value * Present value interest factor at 2% for 14 period

= ($15,255*12.10625) + (508,500*0.75788)

$184,680.84 + $385,382

= $570,063

Issue Price = $570,063

b)

Case B – Issue Price – Market interest rate (annual): 6 percent

Face value of bonds = $508,500

Half yearly maturity period = 7 years * 2 = 14

Semiannual market interest = 6 /2 = 3%

Semiannual Cash Interest = Face value 508,500 * Coupon Rate 6% * ½ = $15,255

Issue Price = Cash Interest * Present value annuity factor at 3% for 14 period + Face value * Present value interest factor at 3% for 14 period

= ($15,255*11.29607) + (508,500*0.66112)

= $172,321.55 + 336,179.52

= $508,501

Issue Price = $508,501

c)

Case C – Issue Price – Market interest rate (annual): 6 percent

Face value of bonds = $508,500

Half yearly maturity period = 7 years * 2 = 14

Semiannual market interest = 8.5 /2 = 4.25%

Semiannual Cash Interest = Face value 508,500 * Coupon Rate 6% * ½ = $15,255

Issue Price = Cash Interest * Present value annuity factor at 4.25% for 14 period + Face value * Present value interest factor at 4.25% for 14 period

= ($15,255*10.39090) + (508,500*0.55839)

= $158,513.18 + 283,941.32

= $442,454

Issue Price = $442,454

Hope the above calculations, working and explanations are clear to you and help you in understanding the concept of question.... please rate my answer...in case any doubt, post a comment and I will try to resolve the doubt ASAP…thank you